### v1

@@ 5.2 Generalization to Diverse Vertex Problems
Firstly, we prove that if $(G,S)\in\mathcal{P}$, then there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that $S=S_{\rho}(\mathcal{D},\alpha)$. Assume that $(G,S)\in\mathcal{P}$. Then there exists an unreduced partial evaluation tree $f$ for $\Phi^{\prime}(z_{1})$ and $G$ such that the root of $f$ is labeled by $S$ and $f$ with pruned external choices evaluates to true. Then for $t_{1},\ldots,t_{n}$ being the leaves of $\mathcal{D}$ there exist partial evaluation trees $f_{1},\ldots,f_{n}$ for $\Phi^{\prime}(z_{1})$ and $G_{t_{1}},\ldots,G_{t_{n}}$ from which we can build $f$ by the induction on the decomposition tree, using the tree product operation in join nodes. For $i\in[n]$, let $S_{i}$ be the label of the root of $f_{i}$. Observe that $S_{i}=\emptyset$ or $S_{i}=\{v\}$ where $t_{i}=Leaf_{\mathcal{D}}(v)$. By the definition of tree product, $S=\bigcup_{i\in[n]}S_{i}$. By the definition of $\mathcal{F}_{t}^{-}$, for each $f_{i}$ there exist $f_{i}^{-}\in\mathcal{F}_{t_{i}}^{-}$ which is isomorphic to it. Observe that two partial evaluation trees $f_{i},f_{i}^{-}$ with roots labeled by ,$\emptyset,\{v\}$ respectively cannot be isomorphic, so the label of the root of $f_{i}^{-}$ is the same as the label of the root of $f_{i}$. By 5.23, there exist exactly one $(\mathfrak{C},\mathcal{D},f^{-})$-witness $\alpha$ such that $\alpha(t_{i})=f_{i}^{-}$ for each $i\in[n]$. Then $S_{\rho}(\mathcal{D},\alpha)=\bigcup_{i\in[n]}S_{i}=S$. Moreover, by Lemma 5.12, $f^{-}$ is isomorphic to $f$. Then $f^{-}$ with pruned external choices evaluates to true, so $\alpha$ is a $(\mathfrak{C},\mathcal{D})$-witness.
@@
Secondly, we prove that if there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that $S=S_{\rho}(\mathcal{D},\alpha)$, then $(G,S)\in\mathcal{P}$. Let $f^{-}=\alpha(t)$ where $t$ is the root of $\mathcal{D}$. Tree $f^{-}$ with pruned external choices evaluates to true. Let $t_{1},\ldots,t_{n}$ be the leaves of $\mathcal{D}$. For $i\in[n]$, let $f_{i}^{-}=\alpha(t_{i})$ and $S_{i}$ be the label of the root of $f_{i}^{-}$. Observe that $S_{i}=\emptyset$ or $S_{i}=\{v\}$ where $t_{i}=Leaf_{\mathcal{D}}(v)$. Then $S_{\rho}(\mathcal{D},\alpha)=\bigcup_{i\in[n]}S_{i}$. Let $f$ be an unreduced partial evaluation tree that can be built from $f_{1}^{-},\ldots,f_{n}^{-}$ by induction on the decomposition tree. Observe that by the definition of tree product, the label of the root of $f$ is $S=\bigcup_{i\in[n]}S_{i}$. By Lemma 5.12, $f$ is isomorphic to $f^{-}$. Then $f$ with pruned external choices also evaluates to true, so $(G,S)\in\mathcal{P}$. ∎
@@
Combining Theorem 5.21 with Theorem 4.6, we get the following corollary, which is a more explicit version of Corollary 1.5 stated in the introduction.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems expressible in $MSO_{1}$, $\Phi_{i}$ be an $MSO_{1}$ formula for $\mathcal{P}_{i}$ and $Div$ be a Venn diversity function. The problem $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time
@@ 6 Conclusions and Open Problems
Our main contribution is that we can convert a monotone DP into one that solves a diverse version of the problem. We think it would be interesting to study this direction past the vertex problems. For example, can we extend the definition of monotonicity to edge problems? In particular, it would be very interesting to study whether there is any hope that we can similarly approach the edge problems. Of course, there is a barrier preventing FPT algorithms for many edge problems, but still, a transformation of the original DP algorithm might exist. Indeed, this is in line with our approach, which also works independently of the running time of the original algorithm (of course, the final running time is dependent on it). Take, for example, the Hamiltonian cycle problem. There is an XP dynamic programming algorithm parameterized by cliquewidth [BKK19]. Is there an algorithm solving the diverse variant of the Hamiltonian cycle problem?
Another intriguing direction would be to research whether $d^{r^{2}}$ overhead in running time comparison of Theorem 1.2 and Theorem 1.3 is really necessary.
We believe that studying the Venn $f$-diversity measures is of independent interest. It would be nice to see some systematic study of differences in such measures or perhaps some experimental study that would give us more understanding of the differences between the various measures. We believe that Venn $f$-diversity measures are rich enough class to study. At this point, let us mention that for our purpose, we can allow even a more general definition of Venn $f$-diversity measure that also has access to the original object, the input graph (as long as we only refer to the final graph $G$ and not to $G_{t}$ for $t$ not being a root of the decomposition). For example, one can design a measure that penalizes/benefits vertices differently based on their degree in $G$. Note that such a definition still allows us to get the same outcome for Theorem 1.2. However, as we currently do not have a good application for this strengthening, we decided to go for a less general definition in the main body for the sake of simplicity and readability.

### v2

@@ 5.2 Dynamic Programming Core for MSO1-expressible Vertex Problems
Firstly, we prove that if $(G,S)\in\mathcal{P}$, then there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that $S=S_{\rho}(\mathcal{D},\alpha)$. Assume that $(G,S)\in\mathcal{P}$. Then there exists an unreduced partial evaluation tree $f$ for $\Phi^{\prime}(z_{1})$ and $G$ such that the root of $f$ is labeled by $S$ and $f$ with pruned external choices evaluates to true. Then for $t_{1},\ldots,t_{n}$ being the leaves of $\mathcal{D}$ there exist partial evaluation trees $f_{1},\ldots,f_{n}$ for $\Phi^{\prime}(z_{1})$ and $G_{t_{1}},\ldots,G_{t_{n}}$ from which we can build $f$ by the induction on the decomposition tree, using the tree product operation in join nodes. For $i\in[n]$, let $S_{i}$ be the label of the root of $f_{i}$. Observe that $S_{i}=\emptyset$ or $S_{i}=\{v\}$ where $t_{i}=Leaf_{\mathcal{D}}(v)$. By the definition of tree product, $S=\bigcup_{i\in[n]}S_{i}$. By the definition of $\mathcal{F}_{t}^{-}$, for each $f_{i}$ there exist $f_{i}^{-}\in\mathcal{F}_{t_{i}}^{-}$ which is isomorphic to it. Observe that two partial evaluation trees $f_{i},f_{i}^{-}$ with roots labeled by $\emptyset,\{v\}$ respectively cannot be isomorphic, so the label of the root of $f_{i}^{-}$ is the same as the label of the root of $f_{i}$. By ˜5.20, there exist exactly one $(\mathfrak{C},\mathcal{D},f^{-})$-witness $\alpha$ such that $\alpha(t_{i})=f_{i}^{-}$ for each $i\in[n]$. Then $S_{\rho}(\mathcal{D},\alpha)=\bigcup_{i\in[n]}S_{i}=S$. Moreover, by Lemma˜5.17, $f^{-}$ is isomorphic to $f$. Then $f^{-}$ with pruned external choices evaluates to true, so $\alpha$ is a $(\mathfrak{C},\mathcal{D})$-witness.
@@
Secondly, we prove that if there exists a $(\mathfrak{C},\mathcal{D})$-witness $\alpha$ such that $S=S_{\rho}(\mathcal{D},\alpha)$, then $(G,S)\in\mathcal{P}$. Let $f^{-}=\alpha(t)$ where $t$ is the root of $\mathcal{D}$. Tree $f^{-}$ with pruned external choices evaluates to true. Let $t_{1},\ldots,t_{n}$ be the leaves of $\mathcal{D}$. For $i\in[n]$, let $f_{i}^{-}=\alpha(t_{i})$ and $S_{i}$ be the label of the root of $f_{i}^{-}$. Observe that $S_{i}=\emptyset$ or $S_{i}=\{v\}$ where $t_{i}=Leaf_{\mathcal{D}}(v)$. Then $S_{\rho}(\mathcal{D},\alpha)=\bigcup_{i\in[n]}S_{i}$. Let $f$ be an unreduced partial evaluation tree that can be built from $f_{1}^{-},\ldots,f_{n}^{-}$ by induction on the decomposition tree. Observe that by the definition of tree product, the label of the root of $f$ is $S=\bigcup_{i\in[n]}S_{i}$. By Lemma˜5.17, $f$ is isomorphic to $f^{-}$. Then $f$ with pruned external choices also evaluates to true, so $(G,S)\in\mathcal{P}$. ∎
˜5.21 concludes the proof of Theorem˜5.18. ∎
@@
Combining Theorem˜5.18 with Theorem˜3.6, we get the following corollary, which is a more explicit version of Corollary˜1.5 stated in the introduction.
@@
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{r}$ be vertex problems expressible in $MSO_{1}$, $\Phi_{i}$ be an $MSO_{1}$ formula for $\mathcal{P}_{i}$, and $Div$ be a Venn diversity measure. The problem $Div^{d}(\mathcal{P}_{1},\ldots,\mathcal{P}_{r})$ on a graph $G$ with given cliquewidth decomposition $\mathcal{D}$ can be solved in time
@@ 6 Conclusions and Open Problems
Our main contribution is that we can convert a monotone DP into one that solves a diverse version of the problem. We think it would be interesting to study this direction past the vertex problems. For example, can we extend the definition of monotonicity to edge problems? However, there is a barrier preventing FPT algorithms for many edge problems, as MSO2 model checking is most likely not even in XP on cliques [CMR00]. Still, a transformation of the original DP algorithm might exist. Indeed, this is in line with our approach, which also works independently of the running time of the original algorithm (of course, the final running time is dependent on it). Take, for example, the Hamiltonian cycle problem. There is an XP dynamic programming algorithm parameterized by the cliquewidth [BKK19]. Is there an XP algorithm solving the diverse variant of the Hamiltonian cycle problem parameterized by the cliquewidth?
Another intriguing direction would be to research whether $d^{r^{2}}$ overhead in running time comparison of Theorem˜1.2 and Theorem˜1.3 is necessary.
We believe that studying the Venn $f$-diversity measures is of independent interest. It would be nice to see some systematic study of differences in such measures or perhaps some experimental study that would give us more understanding of the various measures. Likewise, there has been a study comparing different diversity measures from the perspective of evolutionary computation [WO03].
At this point, let us mention that for our purpose, we can allow even a more general definition of Venn $f$-diversity measure that also has access to the original object. More precisely, we can use properties and shape of the input graph (as long as we only refer to the final graph $G$ and not to $G_{t}$ for $t$ not being a root of the decomposition). For example, one can design a measure that penalizes/benefits vertices differently based on their degree in $G$. Note that such a definition still allows us to get the same outcome for Theorem˜1.2. However, as we currently do not have a good application for this strengthening, we decided to go for a less general definition of the measures in the main body for the sake of simplicity and readability.